OpenStudy (anonymous):

The population of a local species of beetle can be found using an infinite geometric series where a1 = 880 and the common ratio is one forth. Write the sum in sigma notation and calculate the sum (if possible) that will be the upper limit of this population.

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I see immediately that the first term, a, is 880 and that the common ratio, r, is 1/4. Can we use the geometric series sum formula? Yes, because the common ratio, 1/4, is between -1 and 1; in other words, the absolute value of the common ratio is |1/4|<1. Great. Therefore, the sum of this series is \[\sum_{n=1}^{infinity} a*r ^{n-1}\]

OpenStudy (mathmale):

Please substitute 880 for a and 1/4 for r. Calculate this sum.

OpenStudy (anonymous):

220

OpenStudy (anonymous):

is this series divergent

OpenStudy (mathmale):

\[\sum_{n=1}^{infinity} 880(\frac{ 1 }{ 4 }) ^{n-1}=?\]

OpenStudy (mathmale):

Have you a book or online reference handy? We h ave not discussed this topic before. If the condition |r|<1 is met, then the sum in question is equal to \[\frac{ a }{ 1-r }\]

OpenStudy (mathmale):

Again, a=880 and r=1/4. show your stuff!

OpenStudy (anonymous):

880*1/4^0=880

OpenStudy (mathmale):

Have you a book or online reference handy? We h ave not discussed this topic before. If the condition |r|<1 is met, then the sum in question is equal to \[\frac{ a }{ 1-r }\]

OpenStudy (mathmale):

880 is just the first term. a = 880. 1/4 is the common ratio. r=(1/4). thus, substituting these values into \[\frac{ a }{ 1-r }\]

OpenStudy (mathmale):

what do you get? If done correctly, this result will represent the sum of the series.

OpenStudy (mathmale):

Again, if |r|<1, then\[\sum_{n=1}^{infinity}ar ^{n-1}=\frac{ a }{ 1-r }\]

OpenStudy (anonymous):

880/1-1/4

OpenStudy (anonymous):

1173!!!

OpenStudy (mathmale):

Would you please write this down. You'll defnitely see it again later. You're on the right track, but what you have there is open to misinterpretation. Please use the Draw utility: